1. You roll two dice. What is the probability that you roll
(a) doubles?
(b) a sum greater than 10?
(c) doubles or a sum greater than 10?

Answer by stanbon(75887)

You roll two dice. What is the probability that you roll
(a) doubles?
There are 6 doubles.
P(doubles) = 6/36
-----------------------------------
(b) a sum greater than 10?
"10"combos: 6/4;5/5;4/6
"11"combos: 6/5;5/6
"12"combos: 6/6
P(sum greater than 10) = 6/36
-----------------------------------
(c) doubles or a sum greater than 10?
# of (doubles or sum > 10) = 6+6 - 2 = 10
P(doubles or sum > 10) = 10/36
=============================================
Cheers,
Stan H.
==========